You have a right triangle with sides x+2, 2x, and 10(the hypotenuse). Find the value of x.
Answer by stanbon(75887)

You have a right triangle with sides x+2, 2x, and 10(the hypotenuse). Find the value of x.
-----
Use Pythagoras::
(2x)^2 + (x+2)^2 = 10^2
------------------------------
4x^2 + x^2 + 4x + 4 = 100
----
5x^2 + 4x -96 = 0
-----------
x = 4
----
2x = 8
x+2 = 6
-----------------
Cheers,
Stan H.